The market demand for a gallon of Kikkamoo Joy Juice is Q = 1000 – 50P. The owner, Grandma Yocum, wants to produce where the elasticity of demand is unity. What price should she charge and what quantity should be sold to achieve that goal?
Answer #1

Ed= (dQ/dP)*(P/Q)

-1=-50*P/1000 – 50P. (The elasticity is expressed. in negative because of negative relation between price and quantity demanded)

1000 – 50P = 50P (simplification or cross multiplication)

100P = 1000

P = 1000/100 = $10

Q = 1000-50*10 = 500